The Everton College store paid $1564, for an order of 42 calculators. The store paid $10 for each scientific calculator. The others, all graphing calculators, cost the store $54 each. How many of each type of calculators was ordered?
Answer by rfer(16322)

s+g=42 s=42-g
10s+54g=1564
10(42-g)+54g=1564
420-10g+54g=1564
44g=1144
g=26 graphing
42-26=16 scientific